I have a few questions if someone could please help
1)Convert 27% to a fraction?
2)B/y=a Solve for y?
3)B = x Solve for x?
x a
4)r squared=2gh solve for h?
5)(6x10squared) (9x10tothethird) Division problem
(2x10tothezero) (3x10tothefourth)

Answer by checkley71(8403)

27%=.27 TO CHANGE A % TO A DECIMAL JUST MOVE THE DECIMAL 2 PLACES TO THE LEFT & DROP THE % SIGN.
B/Y=A CROSS MULTIPLY GIVES B=YA NOW DIVIDE BOTH SIDES BY A WE GET Y=B/A
B/X=X/A CROSS MULTIPLY WE GET X^2=AB OR X=SQRT(AB)
R^2=2GH OR DIVIDING BOTH SIDES BY 2G WE GET H=R^2/2G
(6*10^2)(9*10^3)/(2*10^0)(3*10^4) OR (600)(9000)/(2)(30000) OR 540000/60000=9